On June 30, 2021, Georgia-Atlantic, Inc. leased warehouse equipment from Builders, Inc. The lease agreement calls for Georgia-Atlantic to make semiannual lease payments of $880,440 over a 3-year lease term (also the asset’s useful life), payable each June 30 and December 31, with the first payment at June 30, 2021. Georgia-Atlantic's incremental borrowing rate is 8.0%, the same rate Builders used to calculate lease payment amounts. Builders manufactured the equipment at a cost of $4.3 million. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
1. Determine the price at which Builders is “selling” the equipment (present value of the lease payments) at June 30, 2021.
2. What amount related to the lease would Builders report in its balance sheet at December 31, 2021 (ignore taxes)?
3. What line item amounts related to the lease would Builders report in its income statement for the year ended December 31, 2021 (ignore taxes)?
(For all requirements, enter your answers in whole dollars and not in millions. Round your final answer to the nearest whole dollar.)
Answer #1

Solution 1:

Semiannual lease payment to be received = $880,440

Total semiannual payments to be received = 3*2 = 6

Incremental borrowing rate = 8%, 4% semiannual

Price at which Builders is “selling” the equipment = Semi Annual lease payments * Cumulative PV Factor of annuity due for 6 periods at 4%

= $880,440 * 5.45182 = $4,800,000

Solution 2:

Semiannual payment received on 30.06.2021 = $880,440

Pretax amount of liability on 30.06.2021 = ($4,800,000 - $880,440) = $3,919,560

Interest revenue for 31.12.2021 = $3,919,560 * 4% = $156,782

Semiannual lease payment received on 31.12.2021 = $880,440

Pre tax amount for receivables on December 31, 2021 = $3,919,560 + $156,782 - $880,440 = $3,195,902

Solution 3:

Pre tax amount of interest Revenue Builder Inc. reports in its income statement = $3,919,560 * 4% = $156,782

Pre tax amount of revenue Builder Inc. reports in its income statement = $4,800,000

Pre tax amount of cost of goods sold to be reported by Builder Inc. in its income statement = $4,300,000

Pretax amounts related to the lease would Builders report in its income statement for the year ended December 31, 2018 = $156,782 + $4,800,000 - $4,300,000 = $656,782